# A Linearly Convergent Frank-Wolfe-type Method for Smooth Convex Minimization over the Spectrahedron

### v1

@@ 2 The Algorithm
Our algorithm is given below as Algorithm 1. As discussed in Section 1.1, our algorithm is inspired by the Away-steps Frank-Wolfe algorithm for optimization over polytopes [16, 21, 12]. Our algorithm applies three different type of steps: I. Standard Frank-Wolfe steps, II. Away/Drop steps, and III. Pairwise steps, all of which are detailed below. In Section 2.1 we discuss in greater detail the efficient implementation of these steps. In particular we show how each step corresponds to a simple leading eigenvector computation. We also describe an implementation that, aside from these eigenvector computations (which can be computed in parallel), only requires $O(n^{2})$ time per iteration.
@@
These are based on moving from the current feasible point towards an extreme point of the feasible set that minimizes the inner-product with the current gradient direction. In case of the spectrahedron, these correspond to a rank-one update of the form ${\mathbf{X}}_{t+1}\leftarrow(1-\eta_{t}){\mathbf{X}}_{t}+\eta_{t}{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$, where ${\mathbf{v}}_{t,+}$ is a unit-length eigenvector corresponding to the leading eigenvalue of $-\nabla{}f({\mathbf{X}}_{t})$ (see also [17, 20, 19]). We set the step size $\eta_{t}\in[0,1]$ using line-search.
@@
Away steps perform rank-one updates in which the weight of some rank-one component supported by the current iterate, is lowered in favor of other components. This takes the from: ${\mathbf{X}}_{t+1}\leftarrow\frac{1}{1-\eta_{t}}\left({{\mathbf{X}}_{t}-\eta_{t}{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right)$, where ${\mathbf{v}}_{t,-}$ is a unit vector in $\textrm{Im}({\mathbf{X}}_{t})$ that maximizes the quadratic product with the current gradient direction. The step size $\eta_{t}$ is set via line search and satisfies $\eta_{t}\in[0,({\mathbf{v}}_{t,-}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{x}}_{t,-})^{-1}]$ in order to maintain the feasibility of ${\mathbf{X}}_{t+1}$ w.r.t. $\mathcal{S}^{n}$ (see Lemma 2 in the sequel). In particular, when $\eta_{t}=({\mathbf{v}}_{t,-}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{x}}_{t,-})^{-1}$ (i.e., the maximal step size allowed), we have that $\textrm{rank}({\mathbf{X}}_{t+1})<\textrm{rank}({\mathbf{X}}_{t})$, and then we refer to this as a drop step. It should be mentioned that this type of away steps closely resembles the ones proposed in [10] for optimization over the matrix nuclear norm ball.
@@
        ${\mathbf{v}}_{t,+}\leftarrow$ leading eigenvector of $-\nabla{}f({\mathbf{X}}_{t})$
        $\eta^{\textrm{FW}}_{t}\leftarrow\textrm{argmin}_{\eta\in[0,1]}f((1-\eta){\mathbf{X}}_{t}+\eta{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top})$
        ${\mathbf{X}}_{t+1}^{\textrm{FW}}\leftarrow(1-\eta^{\textrm{FW}}_{t}){\mathbf{X}}_{t}+\eta^{\textrm{FW}}_{t}{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$
@@
        ${\mathbf{X}}_{t+1}\leftarrow\textrm{argmin}_{{\mathbf{X}}\in\{X_{t+1}^{\textrm{FW}},X_{t+1}^{\textrm{AFW}},X_{t+1}^{\textrm{PFW}}\}}f({\mathbf{X}})$
@@
Let ${\mathbf{X}}\in\mathbb{S}^{n}_{+}$ and ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}})$. Consider the matrix ${\mathbf{Y}}={\mathbf{X}}-\lambda{\mathbf{v}}{\mathbf{v}}^{\top}$ for some $\lambda\geq 0$. If $\lambda\leq({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then ${\mathbf{Y}}\succeq 0$. Moreover, if $\lambda=({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then $\textrm{rank}({\mathbf{Y}})<\textrm{rank}({\mathbf{X}})$ and in particular ${\mathbf{X}}^{\dagger}{\mathbf{v}}\in\textrm{Ker}({\mathbf{Y}})$.
@@ 2.1 Implementation details
${\mathbf{v}}_{t,-}$ could be computed by a standard leading eigenvector computation w.r.t. the matrix ${\mathbf{M}}=\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}+(1+\zeta)\|{\nabla{}f({\mathbf{X}}_{t})}\|\Pi_{t}$, for some arbitrary positive constant $\zeta>0$. Observe that by construction, ${\mathbf{M}}$ is positive semidefinite and its leading eigenvectors, which correspond to a strictly positive leading eigenvalue, indeed must lie in $\textrm{Im}({\mathbf{X}}_{t})$. In particular, ${\mathbf{M}}$ need not be computed explicitly: standard fast iterative leading eigenvector methods (such as Lanczos-type algorithms) rely only on computing matrix-vector products w.r.t. to the matrix ${\mathbf{M}}$, and thus, given $\Pi_{t}$ (or ${\mathbf{X}}_{t}^{\dagger}$), these could be implemented efficiently in $O(n^{2})$ time per such matrix-vector product without explicitly computing ${\mathbf{M}}$.
@@
${\mathbf{u}}_{t,-}$ is a random vector distributed uniformly over the unit sphere in the space $\textrm{Im}({\mathbf{X}}_{t})$. Using the rotation invariance of the multi-variate standard Gaussain distribution $\mathcal{N}(\mathbf{0},{\mathbf{I}}_{n})$, we have that given a random Gaussian vector ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\mathbf{I}}_{n})$, the vector ${\mathbf{u}}_{t,-}=\frac{\Pi_{t}{\mathbf{z}}}{\|{\Pi_{t}{\mathbf{z}}}\|}$ is distributed uniformly over the unit sphere in $\textrm{Im}({\mathbf{X}}_{t})$.
@@
All computations discussed above require maintaining the projection matrix $\Pi_{t}$. Additionally, the computations of the scalars $\lambda_{t},\gamma_{t}$ in Algorithm 1 require also access to the pseudo inverse matrix ${\mathbf{X}}_{t}^{\dagger}$. We now discuss how it is possible to efficiently update either $\Pi_{t}$ of ${\mathbf{X}}_{t}^{\dagger}$ (or both) throughout the run of the algorithm.
@@
Similarly to the well known Sherman-Morrison-Woodbury formula for the fast update of the matrix inverse after a rank-one update, [23] established a formula for updating the pseudo inverse matrix after a rank-one update. While the update in case of the pseudo inverse is substantially more involved, the computational complexity is similar. In particular, on each iteration $t$ of Algorithm 1, given ${\mathbf{X}}_{t}^{\dagger}$ and the quantities ${\mathbf{v}}_{t,-}$, $\lambda_{t}$, ${\mathbf{v}}_{t,+}$, ${\mathbf{u}}_{t,-}$, ${\mathbf{u}}_{t,+}$, $\eta_{t}^{\textrm{FW}}$, $\eta_{t}^{\textrm{AFW}}$, $\eta_{t}^{\textrm{PFW}}$, $\gamma_{t}$, the pseudo inverse matrix of the next iterate, ${\mathbf{X}}_{t+1}^{\dagger}$, can be computed in $O(n^{2})$ time 333while [23] considers the general case of complex non-square matrices which leads to 6 different types of updates for the pseudo inverse, here since we consider symmetric matrices, only 3 of the 6 possibilities need to be considered. For more concrete details see Lemma 9 in the appendix. Since we have that $\Pi_{t+1}={\mathbf{X}}_{t+1}{\mathbf{X}}_{t+1}^{\dagger}$, it is not required to also maintain $\Pi_{t+1}$ explicitly. This leads to the following theorem.
@@
Algorithm 1 admits an implementation such that on each iteration $t$, aside of the computation of the eigenvectors ${\mathbf{v}}_{t,-},{\mathbf{v}}_{t,+},{\mathbf{u}}_{t,+}$, all other computations can be carried out in $O(n^{2})$ time.
@@
A different possibility is not to maintain the pseudo inverse ${\mathbf{X}}_{t}^{\dagger}$ explicitly, but only the projection matrix $\Pi_{t}$. Let us consider the four possible cases. I. In case a drop step is performed, then according to Lemma 2 we have that the unit vector ${\mathbf{z}}={\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-}$ satisfies ${\mathbf{z}}\in\textrm{Im}({\mathbf{X}}_{t})$ and ${\mathbf{z}}\in\textrm{Ker}({\mathbf{X}}_{t+1})$. This implies that $\Pi_{t+1}=\Pi_{t}-\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$. The vector ${\mathbf{z}}$ could be approximated to arbitrary accuracy by solving the simple least squares problem: $\min_{{\mathbf{y}}\in\mathbb{R}^{n}}\|{{\mathbf{X}}_{t}{\mathbf{y}}-{\mathbf{v}}_{t,-}}\|^{2}$. II. In case a standard Frank-Wolfe step is preformed and $\eta_{t}^{\textrm{FW}}<1$, defining ${\mathbf{z}}:={\mathbf{v}}_{t,+}-\Pi_{t}{\mathbf{v}}_{t,+}$, we have that $\Pi_{t+1}=\Pi_{t}+\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$. In case $\eta_{t}^{\textrm{FW}}=1$, we set $\Pi_{t+1}={\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$. III. In case an away step is taken (which is not a drop step), we have that $\textrm{Im}({\mathbf{X}}_{t+1})=\textrm{Im}({\mathbf{X}}_{t})$, and so $\Pi_{t+1}=\Pi_{t}$. IV. If a pairwise step is taken, then we update $\Pi_{t}$ using two steps: a drop step (as in step I above) w.r.t. the direction ${\mathbf{u}}_{t,-}$, and then adding the direction ${\mathbf{u}}_{t,+}$ (as in step II above). That is, denoting ${\mathbf{w}}:={\mathbf{X}}_{t}^{\dagger}{\mathbf{u}}_{t,-}$, $\Pi_{t+1/2}:=\Pi_{t}-\frac{1}{\|{{\mathbf{w}}}\|^{2}}{\mathbf{w}}{\mathbf{w}}^{\top}$, and then denoting ${\mathbf{z}}:={\mathbf{u}}_{t,+}-\Pi_{t+1/2}{\mathbf{u}}_{t,+}$, we have that $\Pi_{t+1}=\Pi_{t+1/2}+\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$. Finally, as with the computation of the product ${\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-}$ in step I, computing the scalars $\lambda_{t},\gamma_{t}$, which requires to compute a matrix-vector product with the pseudo inverse matrix ${\mathbf{X}}_{t}^{\dagger}$, can be done by solving a simple least squares problem.

### v2

@@ 2 The Algorithm
Our algorithm is given below as Algorithm 1. As discussed in Section 1.1, our algorithm is inspired by the Away-steps Frank-Wolfe algorithm for optimization over polytopes [17, 22, 13]. Our algorithm applies three different types of steps: I. Standard Frank-Wolfe steps, II. Away/Drop steps, and III. Pairwise steps, all of which are detailed below. In Section 2.1 we discuss in greater detail the efficient implementation of these steps. In particular we show how each step corresponds to a simple leading eigenvector computation. We also describe an implementation that, aside from these eigenvector computations (which can be computed in parallel), only requires $O(n^{2})$ time per iteration.
@@
These are based on moving from the current feasible point towards an extreme point of the feasible set that minimizes the inner-product with the current gradient direction. In case of the spectrahedron, these correspond to a rank-one update of the form ${\mathbf{X}}_{t+1}\leftarrow(1-\eta_{t}){\mathbf{X}}_{t}+\eta_{t}{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$, where ${\mathbf{v}}_{t,+}$ is a unit-length eigenvector corresponding to the leading eigenvalue of $\nabla{}f({\mathbf{X}}_{t})$ (see also [18, 19, 21]). We set the step size $\eta_{t}\in[0,1]$ using line-search.
@@
Away steps perform rank-one updates in which the weight of some rank-one component supported by the current iterate is lowered in favor of other components. This takes the form: ${\mathbf{X}}_{t+1}\leftarrow\frac{1}{1-\eta_{t}}\left({{\mathbf{X}}_{t}-\eta_{t}{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right)$, where ${\mathbf{v}}_{t,-}$ is a unit vector in $\textrm{Im}({\mathbf{X}}_{t})$ that maximizes the quadratic product with the current gradient direction. The step size $\eta_{t}$ is set via line search and satisfies $\eta_{t}\in[0,({\mathbf{v}}_{t,-}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-})^{-1}]$ in order to maintain the feasibility of ${\mathbf{X}}_{t+1}$ w.r.t. $\mathcal{S}^{n}$ (see Lemma 2 in the sequel). In particular, when $\eta_{t}=({\mathbf{v}}_{t,-}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-})^{-1}$ (i.e., the maximal step size allowed), we have that $\textrm{rank}({\mathbf{X}}_{t+1})<\textrm{rank}({\mathbf{X}}_{t})$, and then we refer to this as a drop step. It should be mentioned that this type of away steps closely resembles the ones proposed in [11] for optimization over the matrix nuclear norm ball.
@@
   ${\mathbf{v}}_{t,+}\leftarrow$ leading eigenvector of $\nabla{}f({\mathbf{X}}_{t})$
   $\eta^{\textrm{FW}}_{t}\leftarrow\textrm{argmin}_{\eta\in[0,1]}f({\mathbf{X}}_{t}+\eta({\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t}))$
   ${\mathbf{X}}_{t+1}^{\textrm{FW}}\leftarrow{\mathbf{X}}_{t}+\eta^{\textrm{FW}}_{t}({\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t})$
@@
   ${\mathbf{X}}_{t+1}\leftarrow\textrm{argmin}_{{\mathbf{X}}\in\{{\mathbf{X}}_{t+1}^{\textrm{FW}},{\mathbf{X}}_{t+1}^{\textrm{AFW}},{\mathbf{X}}_{t+1}^{\textrm{PFW}}\}}f({\mathbf{X}})$
@@
Let ${\mathbf{X}}\in\mathbb{S}^{n}_{+}$ and ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}}),{\mathbf{v}}\neq\mathbf{0}$. Consider the matrix ${\mathbf{Y}}={\mathbf{X}}-\lambda{\mathbf{v}}{\mathbf{v}}^{\top}$ for some $\lambda\geq 0$. If $\lambda\leq({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then ${\mathbf{Y}}\succeq 0$. Moreover, if $\lambda=({\mathbf{v}}^{\top}{\mathbf{X}}^{\dagger}{\mathbf{v}})^{-1}$, then $\textrm{rank}({\mathbf{Y}})=\textrm{rank}({\mathbf{X}})-1$ and in particular ${\mathbf{X}}^{\dagger}{\mathbf{v}}\in\textrm{Ker}({\mathbf{Y}})$.
@@ 2.1 Implementation details
This computation could be reduced to a standard leading eigenvector computation as follows. Computing ${\mathbf{v}}_{t,-}$, as defined in Algorithm 1, corresponds to computing a vector in the set:
$\arg\max\nolimits_{{\mathbf{v}}\in\mathbb{R}^{n}:\Pi_{t}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\nabla{}f({\mathbf{X}}_{t}){\mathbf{v}}$
$=\arg\max\nolimits_{{\mathbf{v}}\in\mathbb{R}^{n}:\Pi_{t}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}{\mathbf{v}}$
$=\arg\max\nolimits_{{\mathbf{v}}\in\mathbb{R}^{n}:\Pi_{t}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\left({\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}+(1+\zeta)\|{\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}}\|\Pi_{t}}\right){\mathbf{v}},$ (5)
where the last equality holds for any real scalar $\zeta$.
For any $\zeta>0$, the matrix ${\mathbf{M}}=\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}+(1+\zeta)\|{\Pi_{t}\nabla{}f({\mathbf{X}}_{t})\Pi_{t}}\|\Pi_{t}$ in the RHS of (2.1) is positive semidefinite, all its eigenvectors with non-zero eigenvalue must lie in $\textrm{Im}(\Pi_{t})=\textrm{Im}({\mathbf{X}}_{t})$, and $\lambda_{1}({\mathbf{M}})>0$. Thus, when solving the maximization problem in the RHS of (2.1), we can drop the explicit constraint $\Pi_{t}{\mathbf{v}}={\mathbf{v}}$, and solve the standard leading eigenvector problem $\max_{{\mathbf{v}}\in\mathbb{R}^{n}:\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}{\mathbf{M}}{\mathbf{v}}$. In particular, $\mathbf{M}$} need not be computed explicitly: standard fast iterative leading eigenvector methods (such as Lanczos-type algorithms) rely only on computing matrix-vector products w.r.t. the matrix $\mathbf{M}$}, and thus, given $\Pi_{t}$ (or ${\mathbf{X}}_{t}^{\dagger}$), these could be implemented efficiently in $O(n^{2})$ time per such matrix-vector product without explicitly computing $\mathbf{M}$}.
@@
${\mathbf{u}}_{t,-}$ is a random vector distributed uniformly over the unit sphere in the space $\textrm{Im}({\mathbf{X}}_{t})$. Using the rotation invariance of the multivariate standard Gaussian distribution $\mathcal{N}(\mathbf{0},{\mathbf{I}}_{n})$, we have that given a random Gaussian vector ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\mathbf{I}}_{n})$, the vector ${\mathbf{u}}_{t,-}=\frac{\Pi_{t}{\mathbf{z}}}{\|{\Pi_{t}{\mathbf{z}}}\|}$ is distributed uniformly over the unit sphere in $\textrm{Im}({\mathbf{X}}_{t})$.
@@
All computations discussed above require maintaining the projection matrix $\Pi_{t}$. Additionally, the computations of the scalars $\lambda_{t},\gamma_{t}$ in Algorithm 1 require also access to the pseudoinverse matrix ${\mathbf{X}}_{t}^{\dagger}$. We now discuss how it is possible to efficiently update either $\Pi_{t}$ or ${\mathbf{X}}_{t}^{\dagger}$ (or both) throughout the run of the algorithm.
@@
Similarly to the well-known Sherman-Morrison-Woodbury formula for the fast update of the matrix inverse after a rank-one update, [24] established a formula for updating the pseudoinverse matrix after a rank-one update. While the update in case of the pseudoinverse is substantially more involved, the computational complexity is similar. In particular, on each iteration $t$ of Algorithm 1, given ${\mathbf{X}}_{t}^{\dagger}$ and the quantities ${\mathbf{v}}_{t,-}$, $\lambda_{t}$, ${\mathbf{v}}_{t,+}$, ${\mathbf{u}}_{t,-}$, ${\mathbf{u}}_{t,+}$, $\eta_{t}^{\textrm{FW}}$, $\eta_{t}^{\textrm{AFW}}$, $\eta_{t}^{\textrm{PFW}}$, $\gamma_{t}$, the pseudoinverse matrix of the next iterate, ${\mathbf{X}}_{t+1}^{\dagger}$, can be computed in $O(n^{2})$ time 333while [24] considers the general case of complex non-square matrices which leads to 6 different types of updates for the pseudoinverse, here since we consider symmetric matrices, only 3 of the 6 possibilities need to be considered. For more concrete details see Lemma 9 in the appendix. Since we have that $\Pi_{t+1}={\mathbf{X}}_{t+1}{\mathbf{X}}_{t+1}^{\dagger}$, it is not required to also maintain $\Pi_{t+1}$ explicitly. This leads to the following theorem.
@@
Algorithm 1 admits an implementation such that on each iteration $t$, aside from the computation of the eigenvectors ${\mathbf{v}}_{t,-},{\mathbf{v}}_{t,+},{\mathbf{u}}_{t,+}$, all other computations can be carried out in $O(n^{2})$ time.
@@
A different possibility is not to maintain the pseudoinverse ${\mathbf{X}}_{t}^{\dagger}$ explicitly, but only the projection matrix onto $\textrm{Im}({\mathbf{X}}_{t})$, $\Pi_{t}$444it may be helpful to recall that for any orthonormal basis for $\textrm{Im}({\mathbf{X}}_{t})$, $\{{\mathbf{q}}_{1},\dots,{\mathbf{q}}_{\textrm{rank}({\mathbf{X}}_{t})}\}$, we have that $\Pi_{t}=\sum_{i=1}^{\textrm{rank}({\mathbf{X}}_{t})}{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\top}$. Let us consider the four possible cases. I. In case a drop step is performed on some iteration $t$, then according to Lemma 2 we have that ${\mathbf{z}}={\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-}$ satisfies ${\mathbf{z}}\in\textrm{Ker}({\mathbf{X}}_{t+1})$. Note also that ${\mathbf{z}}\in\textrm{Im}({\mathbf{X}}_{t})$. This implies that $\Pi_{t+1}=({\mathbf{I}}-\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top})\Pi_{t}=\Pi_{t}-\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$. The vector $\mathbf{z}$} could be approximated to arbitrary accuracy by solving the simple least squares problem: $\min_{{\mathbf{y}}\in\mathbb{R}^{n}}\|{{\mathbf{X}}_{t}{\mathbf{y}}-{\mathbf{v}}_{t,-}}\|^{2}$  [^5] Recall each solution to this LS problem is of the form ${\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-}+{\mathbf{q}}$, for some ${\mathbf{q}}\in\textrm{Ker}({\mathbf{X}}_{t})$. II. In case a standard Frank-Wolfe step is performed and $\eta_{t}^{\textrm{FW}}<1$, defining ${\mathbf{z}}:={\mathbf{v}}_{t,+}-\Pi_{t}{\mathbf{v}}_{t,+}$, we have that $\Pi_{t+1}=\Pi_{t}+\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$ (if ${\mathbf{z}}=\mathbf{0}$, we simply set $\Pi_{t+1}=\Pi_{t}$). In case $\eta_{t}^{\textrm{FW}}=1$, we simply have that $\Pi_{t+1}={\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$. III. In case an away step is taken (which is not a drop step), we have that $\textrm{Im}({\mathbf{X}}_{t+1})=\textrm{Im}({\mathbf{X}}_{t})$, and so $\Pi_{t+1}=\Pi_{t}$. IV. If a pairwise step is taken, then we update $\Pi_{t}$ using two steps: a drop step (as in Case I above) w.r.t. the direction ${\mathbf{u}}_{t,-}$, and then adding the direction ${\mathbf{u}}_{t,+}$ (as in Case II above). That is, denoting ${\mathbf{w}}:={\mathbf{X}}_{t}^{\dagger}{\mathbf{u}}_{t,-}$, $\Pi_{t+1/2}:=\Pi_{t}-\frac{1}{\|{{\mathbf{w}}}\|^{2}}{\mathbf{w}}{\mathbf{w}}^{\top}$, and then denoting ${\mathbf{z}}:={\mathbf{u}}_{t,+}-\Pi_{t+1/2}{\mathbf{u}}_{t,+}$, we have that $\Pi_{t+1}=\Pi_{t+1/2}+\frac{1}{\|{{\mathbf{z}}}\|^{2}}{\mathbf{z}}{\mathbf{z}}^{\top}$ (or simply $\Pi_{t+1}=\Pi_{t+1/2}$ in case ${\mathbf{z}}=\mathbf{0}$). Finally, as with the computation of the product ${\mathbf{X}}_{t}^{\dagger}{\mathbf{v}}_{t,-}$ in Case I, computing the scalars $\lambda_{t},\gamma_{t}$, which requires computing a matrix-vector product with the pseudoinverse matrix ${\mathbf{X}}_{t}^{\dagger}$, can be done by solving a simple least squares problem.
Line-search computations:
There are two such computations in the algorithm, one for computing $\eta_{t}^{\textrm{FW}}$ — the step-size for a standard FW update, and the second for computing $\eta_{t}^{\textrm{AFW}}$ — the step-size for the away step. First, let us see that both of these take the same form. We can rewrite the updates for $\eta_{t}^{\textrm{AFW}}$ and ${\mathbf{X}}_{t+1}^{\textrm{AFW}}$ in Algorithm 1 as follows:
$\eta^{\textrm{AFW}}_{t}$ $\leftarrow\textrm{argmin}_{\eta\in[0,\lambda_{t}]}f\left({{\mathbf{X}}_{t}+\frac{\eta}{1-\eta}\left({{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right)}\right),$
${\mathbf{X}}_{t+1}^{\textrm{AFW}}$ $\leftarrow{\mathbf{X}}_{t}+\frac{\eta^{\textrm{AFW}}_{t}}{1-\eta^{\textrm{AFW}}_{t}}\left({{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right).$
Thus, we can rewrite these updates to have the same form as those for the standard FW update:
$\tilde{\eta}^{\textrm{AFW}}_{t}$ $\leftarrow\textrm{argmin}_{\tilde{\eta}\in[0,\frac{\lambda_{t}}{1-\lambda_{t}}]}f\left({{\mathbf{X}}_{t}+\tilde{\eta}\left({{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right)}\right),$
${\mathbf{X}}_{t+1}^{\textrm{AFW}}$ $\leftarrow{\mathbf{X}}_{t}+\tilde{\eta}^{\textrm{AFW}}_{t}\left({{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\right).$
Such line-search computations are standard in FW algorithms [21, 22]. In particular, when $f$ is quadratic these admit a simple closed-form solution. In other cases, they amount to the minimization of a 1-dimensional convex function over a closed interval which can be done efficiently via bisection. Moreover, since our algorithm already requires the knowledge of the smoothness constant $\beta$, both of these computations could be replaced with performing a line search w.r.t. the local quadratic upper-bound of $f$:
$\eta^{\textrm{FW}}$ $\leftarrow\textrm{argmin}_{\eta\in[0,1]}f({\mathbf{X}}_{t})+\eta\langle{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t},\nabla{}f({\mathbf{X}}_{t})}\rangle+\frac{\eta^{2}\beta}{2}\|{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t}}\|_{F}^{2},$
$\eta^{\textrm{AFW}}$ $\leftarrow\textrm{argmin}_{\eta\in[0,\frac{\lambda_{t}}{1-\lambda_{t}}]}f({\mathbf{X}}_{t})+\eta\langle{{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top},\nabla{}f({\mathbf{X}}_{t})}\rangle+\frac{\eta^{2}\beta}{2}\|{{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top}}\|_{F}^{2}.$
These are quadratic minimization problems and thus admit simple closed-form solutions. In particular, it is a simple observation that using these computations, instead of those in Algorithm 1, will not change our analysis and all theoretical derivations will remain intact.